OpenStudy (anonymous):

Expand the logarithmic expression log(8) a/2

OpenStudy (abb0t):

\[\frac{ \log \frac{ a }{ 2 } }{ \log8 } + a^6\]

OpenStudy (abb0t):

Series expansion @ a = 0

OpenStudy (abb0t):

The expansion is different when a = ∞

OpenStudy (zepp):

I think she just want to write that into two logarithm expression, and not the expansion as we know in calculus... am I right? D:

OpenStudy (anonymous):

right like log(8)2 -log (log8)

OpenStudy (zepp):

Then use this awesome formula :D\[\large \log_c\frac{a}{b}=\log_c a-\log_cb\]

OpenStudy (zepp):

\[\large \log_8 \frac{a}{2}=\log_8?-\log_8?\]

OpenStudy (anonymous):

=log (8)a-log(8)2

OpenStudy (zepp):

Yup, you got it :)